# Bose metals, from prediction to realization

### v1

@@ I Introduction
Quantum phase slips (QPS) are tunnelling events in which the phase on one condensate island makes a flip by an angle ±$\pm 2\pi$ to return in the same position, corresponding to what is known as a large gauge transformation. For some reason, QPS are prominently discussed as fundamental aspects for the low-temperature properties of 1D Josephson junction chains golubev ; arutyunov but are rarely taken into account for JJA and films in 2D. Here too, though they are of paramount importance. Because of the U(1) gauge symmetry, the only gauge invariant degrees of freedom encoded in the grain phases are the vortices living on the plaquettes of the array, i.e. on the dual lattice. Let us consider an infinite line of phase slips of alternating signs ending on a particular site. This is everywhere a pure gauge configuration except on the end site of the line, where the slip corresponds to a vortex tunnelling from one plaquette to an adjacent one, as shown in Fig. 1.
@@
[…]
@@
[…]
@@
[…]
@@
$\left[a_{i}({\bf x}),a_{j}({\bf y})\right]$ $=-i\epsilon^{ij}\delta^{2}\left({\bf x}-{\bf y}\right)\ ,$
$\left[b_{i}({\bf x}),b_{j}({\bf y})\right]$ $=-i\epsilon^{ij}\delta^{2}\left({\bf x}-{\bf y}\right)\ ,$ (9)
@@
[…]
@@
Two points are important to mention about this result. First of all, the edges on which there is metallic charge transport need not be the external edges of the sample; there can be as well internal edges forming a percolation network, exactly as in the quantum Hall states cc . Second, edge charge transport is symmetry-protected from localization and disorder-induced dissipation by its correspondence with a topological bulk ground state bti1 ; bti2 ; bti3 : quantum phase slips are the only dissipation mechanism. It is now established experimentally that vortices are unpaired and mostly frozen in the Bose metal state while a fraction remain mobile and cause the residual metallic sheet resistance saturation marcus3 .
@@
felt by the Cooper pairs, where $N_{\Phi}$ is the number of vortices and $A$ the area of the sample, and the charge distribution with the corresponding uniform “magnetic field”

### v2

@@ I Introduction
Quantum phase slips (QPS) are tunnelling events in which the phase on one condensate island makes a flip by an angle $\pm 2\pi$ to return in the same position, corresponding to what is known as a large gauge transformation. For some reason, QPS are prominently discussed as fundamental aspects for the low-temperature properties of 1D Josephson junction chains golubev ; arutyunov but are rarely taken into account for JJA and films in 2D. Here too, though they are of paramount importance. Because of the U(1) gauge symmetry, the only gauge invariant degrees of freedom encoded in the grain phases are the vortices living on the plaquettes of the array, i.e. on the dual lattice. Let us consider an infinite line of phase slips of alternating signs ending on a particular site. This is everywhere a pure gauge configuration except on the end site of the line, where the slip corresponds to a vortex tunnelling from one plaquette to an adjacent one, as shown in Fig. 1.
@@
$\Phi\left(C_{I},C_{J}\right)$ $={1\over 4\pi}\int_{0}^{1}ds\int_{0}^{1}dt\ {dx_{I}^{\mu}\over ds}\epsilon_{\mu\alpha\nu}{(x_{I}-x_{J})^{\alpha}\over|x_{I}-x_{J}|^{3}}{dx_{J}^{\nu}\over dt}\ .$ (3)
@@
$Q_{\mu}$ $=\sum_{I}Q_{I}\int ds\ {dx_{I}^{\mu}\over ds}\ \delta^{3}\left(x-x_{I}(s)\right)\ ,$
$M_{\mu}$ $=\sum_{J}M_{J}\int dt\ {dx_{J}^{\mu}\over dt}\ \delta^{3}\left(x-x_{J}(t)\right)\ .$ (4)
@@
$S_{\rm top}=\ 2\pi\int d^{3}x\ Q_{\mu}\epsilon_{\mu\alpha\nu}{\partial_{\alpha}\over\nabla^{2}}M_{\nu}\ ,$ (5)
@@
$\left[a_{i}({\bf x}),b_{j}({\bf y})\right]=-i\ 2\pi\ \epsilon^{ij}\delta^{2}\left({\bf x}-{\bf y}\right)\ ,$ (9)
@@
$H_{\rm e}={1\over 2\pi}\int ds\left[v_{b}\left(\partial_{s}\xi\right)^{2}+v_{b}\left(\partial_{s}\eta\right)^{2}\right]\ ,$ (14)
@@
Two points are important to mention about this result. First of all, the edges on which there is metallic charge transport need not be the external edges of the sample; there can be as well internal edges forming a percolation network, exactly as in the quantum Hall states $c$ . Second, edge charge transport is symmetry-protected from localization and disorder-induced dissipation by its correspondence with a topological bulk ground state bti1 ; bti2 ; bti3 : quantum phase slips are the only dissipation mechanism. It is now established experimentally that vortices are unpaired and mostly frozen in the Bose metal state while a fraction remain mobile and cause the residual metallic sheet resistance saturation marcus3 .
@@
felt by the Cooper pairs, where $N_{\Phi}$ is the number of vortices and AA the area of the sample, and the charge distribution with the corresponding uniform “magnetic field”